Find the force necessary to keep the toy car in equilibrium. The mass of the toy is 5kg and the inclination of the plane is 30°. Friction is neglected.


Please show the complete solution.

1
Expert's answer
2021-11-11T12:04:45-0500

The force that is directed along the slope can be found from the fact that the component of the force of gravity 'pulls' the car down the slope. By Newton's second and first laws, equilibrium means that the vector sum of the forces acting on the toy is zero.


"F-mg\sin\theta=0,\\
F=mg\sin30°=24.5\text{ N}."